Find the matrix A for T relative to the basis B. T: R2 → R2, T(x, y) = (5x – y, y - x), B = {(1, -2), (0, 3)} A =
Answer #1

o T. IR² K² - T65,0)=(sx-y,.yaj Basks B= {(1,2), (0,3)} TO 1,-2)= (5+2, -2-1) = (7,-3) = c.lf, -2) +6 (0,3) . C=7, -26, +36=